40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15070     Entry time: Wed Dec 4 08:54:07 2019
Author: Yehonathan 
Type: Update 
Category: IMC 
Subject: Mirror analog shaking 

{Yehonathan, Gavin}

Yesterday we tried to shake ITMX with a function generator in order to observe the 28.8kHz drum mode.

We laid a long BNC cable that runs from the YARM to the XARM. This cable either needs to be collected back to the BNC big plastic cable box under the IMC or be labeled so that it could be found easily in the future.

First, we tried to shake it at a lower frequency (100's of Hz) where the shaking should be easily observed in the POSX channel. We try driving the POS channel on the ITMX servo but nothing happens. Most likely it is disconnected.

While setting up for shaking the individual OSEM channels 4 CDSs crashed (c1lsc, c1ass, c1oaf, c1cal).

 

ELOG V3.1.3-